OpenStudy (anonymous):

How to do this 1 _______ √1-x^4

OpenStudy (campbell_st):

is it \[1/\sqrt{1-x^4}\]

OpenStudy (anonymous):

x^4 has to be less than or equal to 1

OpenStudy (campbell_st):

and what do you have to do..?

OpenStudy (anonymous):

domain maybe?

OpenStudy (campbell_st):

x cannot be 1 or -1

OpenStudy (anonymous):

well the whole problem i will write out

OpenStudy (anonymous):

why can't it be you can have sqrt (0)

OpenStudy (anonymous):

1 is your answer

OpenStudy (anonymous):

well it probably was not √1-4^3

OpenStudy (anonymous):

there was never really a question

OpenStudy (anonymous):

but lemme get the whole problem out first

OpenStudy (anonymous):

oh the 1 is part of the equation..

OpenStudy (campbell_st):

I"ll ray rationalise... \[1/\sqrt{1-x^4} \times \sqrt{1 + x^4}/\sqrt{1 +x^4}\] gives \[\sqrt{1 - x^4}/(1-x^4)\]

OpenStudy (anonymous):

f(x) = x^3 - x^4

OpenStudy (anonymous):

thought he was saying how do you do this one:

OpenStudy (anonymous):

this is the whole problem

OpenStudy (anonymous):

was getting _____ √1-x^4 correct?

OpenStudy (anonymous):

where is that root sign coming from i'm still not sure what you're trying to do

OpenStudy (anonymous):

i just dont know if how i got it was correct entire problem f(x) = x^3 - x^4

OpenStudy (anonymous):

f(x)=x^3(1-x)

OpenStudy (anonymous):

that's just a function you can do ALOT of things with it?

OpenStudy (anonymous):

i am supposed to see if it is even odd or neither

OpenStudy (anonymous):

if it's odd f(x)=-f(-x)

OpenStudy (campbell_st):

well find f(-x) so f(-x) = (-x)^3 - (-x)^4 = -x^3 - x^4 if a function is even f(-x) = f(x) if odd f(-x) = -f(x) in your question the function is neither odd nor even for the above reason

OpenStudy (anonymous):

like he said..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

tyvm ppls

OpenStudy (anonymous):

:)